how do you convert this equation into standard form for an equation of a line that passes through (5,2) and has a y-intercept of 7?
Answer by rfer(16322)

y=mx+b
2=m*5+7
2=5m+7
2-7=5m
-5=5m
-5/5=m
m=-1
---------------
y=-x+7
--------
x+y=7